Hd 0 5 10 15 20 25 30 35 40 45 Volume of NaOH added (ml)

An unknown acid is titrated with a strong base (NaOH) to produce the above titration curve. Key points on this graph reveal important information about the substance and the solution created with it.

  1. What volume of NaOH is needed to reach the equivalence point (the point where you added equal moles of NaOH to the acid)?    Select one of the following : ["10 mL", "5 mL", "20 mL", "30 mL", "40 mL", "25 mL", "15 mL", "35 mL"]      

What is the pKa of this acid?  [ Select ] ["1.8", "3.4", "13.0", "8.0", "1.0", "4.0", "1.6", "2.8"]   

At the equivalence point, is the solution neutral, acidic or basic. Why?    [ Select ]    ["neutral, all the acid and base have neutralized each other", "acidic, acids produce acidic solutions", "basic, the conjugate of a weak acid is a weak base", ""]      
Answer #1

The halfway of steep curve is the equivalence point.

The volume of NaOH at half equivalence point = 30 mL.

so, at half equivalence point volume of NaOH will be

= (30/2) = 15 mL.

pH at half equivalence point = 2.8

Before equivalence point solution acts as buffer

At half equivalence point

[HA[ = [NaA].

So, by Henderson-Hassebalch equation is

pH = pKa + log [NaA]/[HA]

So, pH = pKa = 2.8

From, the curve pH at equivalence point = 8.

at equivalence point solution is basic ( the conjugate of a weak acid is a weak base ).

y half equivalence point equivalence point 21 - 15 . 30 vol of Naoh